# Superconductivity in Epitaxial SiGe for Cryogenic Electronics

### v1

@@ III Conclusions
In conclusion, our results establish that MBE growth of hyperdoped Ga:Ge exhibits superconductivity arising from substitutional Ga incorporation, allowing for the creation of coherent semiconductor-superconductor interfaces. TEM and EDS imaging confirm pristine structural and compositional homogeneity. Synchrotron-based methods confirm Ga occupies substitutional sites within the Ge lattice and the interplay between Ga incorporation and epitaxial constraints (substrate clamping) leads to a subtle tetragonal (-Sn) distortion which aligns with our DFT structural model. Electronic structure calculations show that Ga incorporation at substitutional sites shifts the Fermi level into the valence band and flattens electronic bands at the R point, likely promoting superconductivity. Our findings highlight the intrinsic nature of superconductivity in hyperdoped Ga:Ge and underscore its potential as a scalable platform for epitaxial semiconductor-superconductor heterostructures using the mature group IV semiconductor toolkit.
IV Supplementary information
S4.1 Methods
@@ IV Supplementary information
To calculate the electronic band structure and density of states, we adopted the MBJLDA meta-GGA functional [41] based on the modified Becke-Johnson exchange functional. This functional has been shown to yield excellent results for the band structure and band gap of many semiconductors, including Ge [42, 43], which standard semilocal DFT predicts erroneously to be metallic. Spin-orbit coupling was included in all electronic structure calculations. A denser k-mesh was employed to compute the density of states. To accurately determine the Fermi-level shifts in Ga-doped Ge, a k-mesh was used.
Scanning transmission electron microscopy: Cross-sectional specimens for scanning transmission electron microscopy (S/TEM) analysis were prepared using a focused ion beam (FIB) system (FEI Helios NanoLab 600 DualBeam). To minimize surface damage, the samples were thinned and polished with Ga ions at 5 kV. S/TEM imaging and energy-dispersive X-ray spectroscopy (EDS) were conducted using a Thermo Fisher Scientific Themis-Z microscope. The instrument operated at an accelerating voltage of 300 kV with a semi-convergence angle of 20 mrad. HAADF-STEM images were collected using a 64–200 mrad high-angle annular dark field (HAADF) detector. High-resolution images were acquired through multiple rapid scans (2048 2048 pixels, 200 ns per pixel), which were combined to improve the signal-to-noise ratio. EDS analysis was performed with the Super-X EDS detector, and the elemental mappings are presented as net count images.
Raman Spectroscopy: Raman spectra were recorded from the thin film surface in a backscattering configuration, using an integrated Edinburgh confocal microRaman instrument (RM5). Excitation was provided by one of three solid-state laser lines (532 nm, 638 nm and 785 nm) and focused on the surface using a long working lens (Olympus, 10×, 0.4 NA). Dispersion was achieved through a 1200 g/mm diffraction grating (spectral resolution 0.2 ), instrument calibration was verified through checking the position of the Si band at 520.7 , and spectra were recorded using a thermoelectric CCD detector. Changes to the local vibrational structure of the hyperdoped film are first confirmed through the evaluation of the excitation energy dependence of the Raman back-scattering signals (Figure S7). At longer excitation wavelengths (785 nm), both the control and target films exhibit comparable Raman response, while more surface-sensitive information is recovered at shorter wavelengths (532 nm) to resolve differences in the thin films (Figure 3f).
@@
EXAFS fine-structure modeling: Fits was conducted in the ARTEMIS, part of the IFEFFIT software package[48]. The scattering paths used to evaluate and model the data were derived from the optimized crystal structures with different defects. Using these input structures to derive the scattering paths, the only agreeable fine structure accounting for the dominant EXAFS signal arising from an R fitting window of 1.8 to 5.1 Å emerged using the calculated Ga substituted structure. To simulate the Ga substituted Ge crystal fine structure, the input paths out to 5 Å are determined using the DFT-derived Ga substitutionally doped Ge structure and the FEFF package[48], with a distance fuzz = 0.030 Å. This generated 6 unique scattering pairs centered on the Ga substitutional site, with some other lower ranked paths omitted from the actual model to avoid over-fitting (see Table S2). A common scaling factor has been used on all path distance corrections to account for the difference between the DFT simulated unit cell volume and the experiment crystal volume at room temperature. A phase correction has been made to Figure 2d in the main text using the most intense Ga-Ge single scattering path, near 2.45 Å. The data were ultimately fit using 18 variables and 26 independent points, resulting in a reasonable R-factor of 0.015. The fitting was conducted in R-space using multiple k-weightings, with k = 1, 2 and 3. Errors of individual fit parameters were determined using ARTEMIS to take into account the correlations between parameters and known parameters without error estimates were fixed during the fit.
@@
S4.2 Supporting Data
| Structure | Lattice Parameter (Å3) | Volume (Å3) | Crystal Structure | (eV) |
@@
Acknowledgements.
NYU team acknowledges funding support from the United States Air Force Office of Scientific Research award number FA9550-21-1-0338. We acknowledge computational resources at the NCI National Facility through the National Computational Merit Allocation Scheme. J.A.S. acknowledges financial support from the Australian Research Council (DE230100173); Y-H.C. and P.J. acknowledge support from the Australian Research Council (LP210200636, CE170100009); C.V. acknowledges financial support from the Australian Research Council (DE220101147). This research was undertaken on the X-ray Absorption Spectroscopy and the SAXS/WAXS beamlines at the Australian Synchrotron, part of ANSTO. The XRR data reported in this paper were obtained at the Central Analytical Research Facility operated by Research Infrastructure (QUT). Electron microscopy was performed at the Center for Electron Microscopy and Analysis (CEMAS) at The Ohio State University. S.S.-R. acknowledges Dan Huber for assisting with TEM sample preparation.References

### v2

@@ Abstract
In conclusion, our results establish that MBE growth of hyperdoped Ga:Ge exhibits superconductivity arising from substitutional Ga incorporation, allowing for the creation of coherent semiconductor-superconductor interfaces. TEM and EDS imaging confirm pristine structural and compositional homogeneity. Synchrotron-based methods confirm Ga occupies substitutional sites within the Ge lattice and the interplay between Ga incorporation and epitaxial constraints (substrate clamping) leads to a subtle tetragonal (-Sn) distortion, in line with our DFT results. Electronic structure calculations show that Ga incorporation at substitutional sites shifts the Fermi level deep into the valence band and flattens the electronic bands near the R point, likely promoting superconductivity. Our findings highlight the intrinsic nature of superconductivity in hyperdoped Ga:Ge and underscore its potential as a scalable platform for epitaxial semiconductor-superconductor heterostructures using the mature group IV semiconductor toolkit.
Acknowledgments
S.S.-R. acknowledges Dan Huber for assisting with TEM sample preparation. P.J. acknowledges fruitful discussions with Ross McKenzie and Ben Powell.
Funding:
P.J.S and J.S. acknowledge funding support from the United States Air Force Office of Scientific Research award number FA9550-21-1-0338. We acknowledge computational resources at the NCI National Facility through the National Computational Merit Allocation Scheme. This work was partially supported by the Australian Research Council under the following grants: DE230100173 (J.A.S.), LP210200636 (Y.-H.C., P.J.), DE220101147 (C.V.). This research was undertaken on the X-ray Absorption Spectroscopy and the SAXS/WAXS beamlines at the Australian Synchrotron, part of ANSTO. The XRR data reported in this paper were obtained at the Central Analytical Research Facility operated by Research Infrastructure (QUT). Electron microscopy was performed at the Center for Electron Microscopy and Analysis (CEMAS) at The Ohio State University.
Author contributions:
Conceptualization: P.J., J.S., and E.D. Methodology: J.A.S., P.J.S., C.V., S.S.-R. Investigation: J.A.S., P.J.S., C.V., A.B., A.D., Y.-H.C., J.V.D., S.S.-R. Visualization: J.A.S., P.J.S. Funding acquisition: J.A.S., P.J., J.S. Project administration: J.A.S., P.J.S., P.J., J.S. Supervision: J.A.S., P.J., J.S., E.D., L.W. Writing – original draft: P.J., J.A.S., P.J.S. Writing – review and editing: J.A.S., P.J.S., C.V., P.J., J.S.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Raw data available on Zenodo: doi.org/10.5281/zenodo.14782527
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S14
Tables S1 to S2
References (23-0)
Supplementary Materials for
Coherent Superconductor-Semiconductor Epitaxy for Integrated Quantum Electronics
Julian A. Steele∗†,
Patrick J. Strohbeen†,
Carla Verdi,
Ardeshir Baktash,
Alisa Danilenko,
Yi-Hsun Chen,
Jechiel van Dijk,
Lianzhou Wang,
Eugene Demler,
Salva Salmani-Rezaie,
Peter Jacobson,
Javad Shabani∗
∗Corresponding author. Email: julian.steele@uq.edu.au, jshabani@nyu.edu
†These authors contributed equally to this work.
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S14
Tables S1 to S2
Materials and Methods
@@ Acknowledgments
To calculate the electronic band structure and density of states, we adopted the MBJLDA meta-GGA functional [29] based on the modified Becke-Johnson exchange functional. This functional has been shown to yield excellent results for the band structure and band gap of many semiconductors, including Ge [30, 31], which standard semilocal DFT predicts erroneously to be metallic. Spin-orbit coupling was included in all electronic structure calculations. A denser k-mesh was employed to compute the density of states. To accurately determine the Fermi-level shifts in Ga-doped Ge, a k-mesh was used. Vibrational properties were computed using finite differences [32] within supercells containing 216 atoms ( supercells of the 8-atom unit cell), employing -point meshes and displacement amplitudes of 0.015 Å.
Scanning transmission electron microscopy: Cross-sectional specimens for scanning transmission electron microscopy (S/TEM) analysis were prepared using a focused ion beam (FIB) system (FEI Helios NanoLab 600 DualBeam). To minimize surface damage, the samples were thinned and polished with Ga ions at 5 kV. S/TEM imaging and energy-dispersive X-ray spectroscopy (EDS) were conducted using a Thermo Fisher Scientific Themis-Z microscope. The instrument operated at an accelerating voltage of 300 kV with a semi-convergence angle of 20 mrad. HAADF-STEM images were collected using a 64-200 mrad high-angle annular dark field (HAADF) detector. High-resolution images were acquired through multiple rapid scans (2048 2048 pixels, 200 ns per pixel), which were combined to improve the signal-to-noise ratio. EDS analysis was performed with the Super-X EDS detector, and the elemental mappings are presented as net count images.
Raman Spectroscopy: Raman spectra were recorded from the thin film surface in a backscattering configuration, using an integrated Edinburgh confocal micro-Raman instrument (RM5). Excitation was provided by one of three solid-state laser lines (532 nm, 638 nm and 785 nm) and focused on the surface using a long working lens (Olympus, 10x, 0.4 NA). Dispersion was achieved through a 1200 g/mm diffraction grating (spectral resolution 0.2 cm-1), instrument calibration was verified through checking the position of the Si band at 520.7 cm-1, and spectra were recorded using a thermoelectric CCD detector. Changes to the local vibrational structure of the hyperdoped film are first confirmed through the evaluation of the excitation energy dependence of the Raman back-scattering signals (Figure S10). At longer excitation wavelengths (785 nm), both the control and target films exhibit comparable Raman response, while more surface-sensitive information is recovered at shorter wavelengths (532 nm) to resolve differences in the thin films (Figure 3f).
@@
EXAFS fine-structure modeling: Fits was conducted in the ARTEMIS, part of the IFEFFIT software package [37]. The scattering paths used to evaluate and model the data were derived from the optimized crystal structures with different defects. Using these input structures to derive the scattering paths, the only agreeable fine structure accounting for the dominant EXAFS signal arising from an R fitting window of 1.8 to 5.1 Å emerged using the calculated Ga substituted structure. To simulate the Ga substituted Ge crystal fine structure, the input paths out to 5 Å are determined using the DFT-derived Ga substitutionally doped Ge structure and the FEFF package [37], with a distance fuzz = 0.030 Å. This generated 6 unique scattering pairs centered on the Ga substitutional site, with some other lower ranked paths omitted from the actual model to avoid over-fitting (see Table LABEL:exafs-table). A common scaling factor has been used on all path distance corrections to account for the difference between the DFT simulated unit cell volume and the experiment crystal volume at room temperature. A phase correction has been made to Figure 2d in the main text using the most intense Ga-Ge single scattering path, near 2.45 Å. The data were ultimately fit using 18 variables and 26 independent points, resulting in a reasonable R-factor of 0.015. The fitting was conducted in R-space using multiple k-weightings, with k = 1, 2 and 3. Errors of individual fit parameters were determined using ARTEMIS to take into account the correlations between parameters and known parameters without error estimates were fixed during the fit.
@@
| Structure | Lattice Parameter (Å) | Volume (Å3) | Crystal Structure | (eV) |